OpenStudy (anonymous):

An inheritance of $800,000 is to be divided among Avery, Connor, and Olivia in the following manner: Olivia is to receive 3/4 of what Connor gets, while Avery gets 1/4 of what Connor gets. How much does each receive?

OpenStudy (anonymous):

Formulate some equations: A+C+O=80000 O=3/4C A=1/4C Solve by substutution.

OpenStudy (anonymous):

so would the equation set up look like this 1/4C+C+3/4C=800000 and then i solve for C and continue from there?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thank you very much! i was able to solve it and got the correct answer :)

OpenStudy (anonymous):

Glad to help

OpenStudy (anonymous):

:)